OpenStudy (anonymous):

A study shows that 5% of airplane flights arrived before the scheduled arrival time and 2.3% arrived more than 30 minutes after the schedule time. a) Calculate P (|T| <= 5) b) On a given day, 500 flights are due to arrive at these airports. What is the probability that at least 50 planes arrive within 5 minutes of the schedule time?

OpenStudy (kirbykirby):

what is the random variable T representing here? Is there a distribution you're supposed to assume for T ?

OpenStudy (anonymous):

@kirbykirby I left the first question out because I thought the info was irrelevant, and since I solved it, I didn't need help for that, but I'll post it here for clarity! a) Let T be the random variable 'difference between scheduled time and actual arrival time' of the flights at these airports. Assuming that T follows a normal distribution, determine the value of its parameters.

OpenStudy (kirbykirby):

It's always safer to post all the necessary info, because it may be relevant :) Regardless, you say you have solved it anyway?

OpenStudy (kirbykirby):

or is it you need help with b) only?

OpenStudy (anonymous):

I need help with a and b from the first post please :)

OpenStudy (anonymous):

I don't understand how to calculate P(|T| <=5)

OpenStudy (kirbykirby):

|T| <= 5 is the same as -5 <= T <= 5 So find P(-5 <= T <= 5)

OpenStudy (kirbykirby):

And the problem gives you the information that P(T > 30) = 0.023

OpenStudy (anonymous):

ok, I got all the way to (-5 <= T <= 5) before getting stuck. I understand that I'm supposed to look for T in the above equation right?

OpenStudy (kirbykirby):

actually I have to go sorry :( .. if no one helps in the meantime I will get back with you later.

OpenStudy (anonymous):

ok sure! thanks though!

OpenStudy (anonymous):

@tkhunny @ganeshie8 are you guys able to assist me? :x

OpenStudy (anonymous):

If a plane lands as scheduled, then \(T=0\), so before the scheduled time you'd have \(T<0\) and after you'd have \(T>0\). You're given that \(P(T<0)=0.05\) and \(P(T>30)=0.023\). You can use this info to find the mean time difference and standard deviation using the standard normal distribution. \[\frac{T-\mu}{\sigma}=Z\] \[\begin{align*} P(T<0)&=P(\mu+Z\sigma<0)\\ 0.05&=P\left(Z<-\frac{\mu}{\sigma}\right) \end{align*}~~~~~~~~~~\Rightarrow~~(1)~~~~-\frac{\mu}{\sigma}=1.65\] \[\begin{align*} P(T>30)&=P(\mu+Z\sigma>30)\\ 0.023&=P\left(Z<\frac{30-\mu}{\sigma}\right) \end{align*}~~~~\Rightarrow~~(2)~~~~\frac{30-\mu}{\sigma}=2.83\] Solve for the mean \(\mu\) and standard deviation \(\sigma\). Or do you already have this part done?

OpenStudy (anonymous):

Actually, (2) should be about \[\frac{30-\mu}{\sigma}=1.99\]

OpenStudy (anonymous):

And (1) should be \[-\frac{\mu}{\sigma}=-1.65\] Sorry about that!

OpenStudy (anonymous):

Once you have the mean/std dev, you can easily compute the probability in part (b). Like @kirbykirby said, \[\begin{align*}P(|T|\le5)&=P(-5\le T\le5)\\ &=P(T\le5)-P(T\le-5)\\ &=P\left(\frac{T-\mu}{\sigma}\le\frac{5-\mu}{\sigma}\right)-P\left(\frac{T-\mu}{\sigma}\le-\frac{5-\mu}{\sigma}\right)\\ &=P\left(Z\le\frac{5-\mu}{\sigma}\right)-P\left(Z\le-\frac{5-\mu}{\sigma}\right) \end{align*}\] You can simplify this a bit using symmetry, but I think the computation is pretty straightforward from here.

OpenStudy (anonymous):

@SithsAndGiggles thanks! I already have the first part down, and I now understand the second part. I have trouble with the last question though D: (b) On a given day, 500 flights are due to arrive at these airports. What is the probability that at least 50 planes arrive within 5 minutes of the schedule time?)

OpenStudy (anonymous):

The answer you get for \(P(|T|<5)\) is the proportion of planes that land within 5 minutes. Treat this as a binomial success probability. If \(X\) is the random variable for number of planes landing within 5 minutes, then \(p=P(|T|<5)\). The probability that at least 50 of 500 planes will land within 5 minutes is \[\large P(X\ge50)=\sum_{k=50}^{500}\binom{500}{k}p^{k}(1-p)^{500-k}\]

OpenStudy (anonymous):

how did you know/ decide to use the binomial success probability? am I correct to assume then, that binomcdf is used for this question? and since we are looking for P(X=>50), it is also written as 1-P(X<=49)?

OpenStudy (anonymous):

Yes, \(P(X\ge50)=1-P(X\le49)\). If that makes computation easier for you, then by all means. The reason for using a binomial variable is that we're concerned with a particular set of planes that land. We're also not worried about the difference in scheduled and actual arrival time. We have a probability associated with the planes of interest (those that land within 5 minutes of schedule), but we don't have anything that suggests how many planes meet this condition, so we have to introduce a discrete random variable. The main clue for using a binomial distribution was the fact that we're concerned with planes satisfying \(|T|<5\) as opposed to \(|T|>5\). Right away we can establish a success/fail relationship between two types of planes. Does this make sense? The reasoning is kind of second nature to me, but actually explaining it is hard to do...

OpenStudy (anonymous):

oh wow! thank you so much! I have been trying to review my probabilities unit and binomial, poisson, normal etc and squeezing them all in 1.5 weeks has been confusing me as they all have the cdf/pdf components and I get confused which ones I should use when. so planes who are late 'fail' and planes who arrive on time/early = 'success' in the success/fail relationship?

OpenStudy (anonymous):

Almost! Planes that "fail" land either more than 5 minutes before or after they're expected, meaning \(T<-5\) or \(T>5\). Planes that succeed land within the interval, \(-5<T<5\).

OpenStudy (anonymous):

ok. so if I were to plug everything into my GDC in binoncdf, the will the trial value be 50 or 500? I know the probability is 0.137 (found from before, and the X value will be 49.

OpenStudy (anonymous):

oh nevermind; the value would be 500 because that is the sample size?

OpenStudy (anonymous):

The parameters of a binomial distribution are the size of the population (in this case, \(n=500\)) and the success probability (\(p=P(|T|<5)\)) Merely arriving on time translates to \(T=0\), which is a point. For continuous distributions, the probability of obtaining a point is zero, \(P(T=k)=0\). The takeaway is that you have to be careful about what you mean by "on time." I've never actually used binomcdf and all those functions on a calculator (rather, I use the pdf and cdf formulas themselves). The total number of trials 500 though.

OpenStudy (anonymous):

ok thank you so much! I hope you'll still be online to answer more (of my) questions :P

OpenStudy (anonymous):

You're welcome! I'll be around...